OpenStudy (anonymous):

Tabitha wants to hang a painting in a gallery but the painting and frame must have an area of 58 square feet. The painting is 7 feet wide by 8 feet long. The frame must have an equal width of x on each side. Which quadratic equation represents the area of the painting and frame combined

OpenStudy (anonymous):

58 = x² Since the painting is on the frame, you can forget the area taken by the painting. Since x>8, because the painting is 8 by 7, there is no solution. Therefore, she can't hang the painting in that particualr gallery.

OpenStudy (anonymous):

forgot choices :/ x2 + 15x − 2 = 0 x2 + 15x + 58 = 0 4x2 + 30x − 2 = 0 4x2 + 30x + 58 = 0